I need answer on number 7 please
5. Students in a statistics class took their first test. The following table lists the scores they earned.
67
67
76
47
85
70
87
76
67
72
84
98
84
64
65
82
81
81
88
74
87
83
Complete the following frequency distribution table using 6 classes: 40-49, 50-59, 60-69, 70-79, 80-89, and 90-99.
Scores

Tally

Frequency

Relative Frequency

Cumulative Relative Frequency

40 - 49





50 - 59





60 - 69





70 - 79





80 - 89





90 - 99





6. Answer the following questions based on the test score data given in Question #5:
a. What is the range of the test scores?
b. What is the median of the test scores?
c. What is the mode of the test scores?
7. Answer the following questions based on the frequency distribution table you got in Question #5:
a. What percentage of the test scores was at least 60?
b. What percentage of the test scores was at most 89?
Answer #1

a. percentage of the test scores was at least 60=0.9545*100=95.45% (See last column "Cumulative relative frequency(>type)" third entry)

b. percentage of the test scores was at most 89=0.9546*100=95.46% (See 2nd last column "Cumulative relative frequency(<type)" 5th entry).

However these two answers are same i.e. 21/22*100=95.45% and the difference arises due to rounding off upto 4 decimal places.